Question: How many ways can 3 students be placed in a row of 5 desk?
I tried the standard way, like if you had 5 students in 5 desk, you do 5x4x3x2x1, but I didn't get how to do this problem.

Answer by Maths68(1474)

First Student can be placed = 5 ways
1 desk is occupied 2nd student can be placed on 4 remaing desks = 4 ways
2 desks are occupied 3rd student can be placed on 3 remaing desks = 3 ways
Total Number of ways = 5 * 4 * 3
Total Number of ways = 60

Or

Use the formula
nPr = n!/(n-r)!
Here we have
n=5
r=3
5P3=5!/(5-3)!
5P3=5!/(2)!
5P3=5*4*3*2*1/2*1
5P3=120/2
5P3=60 ways.